# Limitations of Data-Driven Spectral Reconstruction – Optics-Aware Analysis and Mitigation

### v2

@@ Abstract.
ccs: Computing methodologies Hyperspectral imagingccs: Computing methodologies Computational photography
Figure 1. Spectral image formation models used in the analysis in this work. Top: In the NTIRE spectral recovery challenges, an RGB image is considered as a linear projection from a high-dimensional hyperspectral datacube to a 3D color image. The existence of metamerism results in identical RGB images for different spectra. The neural network trained in this way cannot distinguish their corresponding spectra. Bottom: A possible mitigation to this problem is to include the optical aberrations of the lens in the image formation model. Spectral information is encoded into the aberrated RGB images, enabling the neural network to tell the difference between metamers. In both cases, the RGB image differences are shown on the right (intensity enhanced for better visualization).
1. Introduction
@@
(Hyper)spectral imaging is a method that involves recording the light in a scene in the form of many, relatively narrow, spectral bands, rather than projected into three broadband RGB color channels. Where RGB imaging utilizes the trichromaticity theory of human color vision, spectral imaging provides additional information that can help discriminate between different materials and lighting conditions that are hard to tell apart in RGB images. For example, red stains in a crime scene could be blood, or paint, or a dyed cloth, which cannot be distinguished from their RGB colors. Skin tumors could not be diagnosed from surrounding tissues of the same color. It is difficult to spot and sort out plastic leaves from living plants by their greenish colors. Therefore, spectral imaging has been applied in many fields, including computer graphics (Kim, 2013), machine vision (Fauvel et al., 2012; Rangnekar et al., 2022), healthcare (Lu and Fei, 2014), agriculture (Dale et al., 2013), and environment (Banerjee et al., 2020), to name just a few.
@@
However, conventional hyperspectral cameras require scanning mechanisms (Lodhi et al., 2019; Huang et al., 2022) to acquire the 3D hyperspectral datacube with 2D sensors. To simplify the demanding hardware, extensive efforts have been made in the development of various snapshot hyperspectral cameras (Wagadarikar et al., 2008; Saragadam and Sankaranarayanan, 2019; Jeon et al., 2019).
@@
On the extreme end of these hardware simplification efforts, deep learning methods have emerged in recent years that attempt to solve the problem entirely in software by reconstructing spectral data from RGB images. This has resulted in three CVPR-hosted NTIRE challenges (Arad et al., 2018, 2020, 2022) and various network architectures (Shi et al., 2018; Li et al., 2020; Zhao et al., 2020; Cai et al., 2022b; Zhang et al., 2022). Yet it remains unclear how these methods generalize to unseen data, how they deal with the difficult but important problem of resolving metamerism (Hill, 1999; Ortega et al., 2020), and how they depend on the optical system of both the RGB source and the spectral cameras used to capture the datasets.
@@
At its core, estimating spectral information from RGB colors is an ill-conditioned one-to-many mapping problem. As stated By Pharr et al. (Pharr et al., 2023) (Ch. 4), “any such conversion is inherently ambiguous due to the existence of metamers”. Intuition would therefore suggest that the achievable spectral fidelity of RGB to spectral methods is limited. However, this intuition flies in the face of very high numerical test results reported in recent NTIRE challenges (Arad et al., 2018, 2020, 2022).
@@
One possible explanation for the experimental success of spectral reconstruction from an RGB image is that the networks learn to exploit spatial structure to reconstruct spectral information. However, in practice metameric or near-metameric colors often occur in situations where the spatial structure is also similar, e.g. vein finding (the global geometry is a segment of a forearm, but vein structure is unknown), or biometrics (e.g., distinguishing real faces from masks or images). To illustrate this effect, we show an example from the smaller and older CAVE dataset (Yasuma et al., 2010) fake_and_real_food_ms, where several fake and real objects are presented in two groups, as shown in Fig. 2. Two sample points from the two red peppers (one real and the other fake, but it’s unclear which is which) have the RGB values [86, 21, 10] and [86, 21, 8], respectively. As can be seen, the real spectra of the real and fake red peppers differ substantially in the red part of the spectrum. A pre-trained model (arbitrarily chosen from previous NTIRE challenges trained on the much larger and newer ARAD1K dataset (Arad et al., 2022)) predicts spectra that differ from the ground truth, but more importantly, the predicted spectra for the real and fake pepper are almost identical, illustrating the failure to adequately deal with metamers.
@@
Figure 2. An example scene fake_and_real_food_ms from the CAVE dataset (Yasuma et al., 2010) consists of objects with visually similar colors, but actually different spectra. Left: Color image with highlighted points on the red peppers. Their RGB values are nearly the same. Right: Ground-truth and reconstructed spectra at the corresponding points show their spectral differences. The reconstructed spectra are predicted by the pre-trained MST++ model (Cai et al., 2022b) on the ARAD1K dataset (Arad et al., 2022). The neural network struggles to distinguish either the two spectra from each other, or from their true spectra.
@@
  • Crucially, the limitations of the datasets that we document not only affect the RGB to spectral work, but also any other spectral reconstruction and processing that uses the same training data (Jacome et al., 2023; Xu et al., 2024; Arguello et al., 2021). We show how metameric augmentation can be used to at least partially overcome the dataset issues.
@@
Through the evidence in this work, we contribute to a deepened understanding of the limitations of current datasets as well as of underlying sources that result in the limitations of spectral reconstruction accuracy. The results of the interplay between metameric spectra and optical aberrations opens the door for new approaches for spectral recovery down the road.
2. Related Work
2.1. Hyperspectral cameras
@@
Conventional hyperspectral imaging systems require filter wheels, liquid-crystal tunable filters, or mechanical motion (e.g., pushbroom) (Lodhi et al., 2019; Huang et al., 2022) to scan the 3D hyperspectral datacube. To enable snapshot acquisition, coded-aperture snapshot spectral imager (CASSI) (Wagadarikar et al., 2008; Choi et al., 2017b) has been proposed to achieve high spectral accuracy using spectrum-dependent coded patterns. Recent methods also exploit spectrally encoded point spread functions (PSFs) to computationally reconstruct a hyperspectral image (Jeon et al., 2019; Baek et al., 2017; Cao et al., 2011). In general, great efforts have been made to simplify hyperspectral camera hardware by software reconstruction.
2.2. Spectral reconstruction from RGB images
@@
A recent trend to solve the snapshot hyperspectral imaging problem is to exploit hyperspectral data with deep neural networks to reconstruct spectral information from RGB images (Arad and Ben-Shahar, 2016). Owing to the wide availability of RGB cameras, this approach seems to be a promising candidate for hyperspectral imaging if successful. A large number of neural network architectures have been proposed in the past three NTIRE spectral recovery challenges  (Arad et al., 2018, 2020, 2022). Our analysis in this paper focuses on this class of methods to gain insights on their strengths and limitations.
2.3. Dataset bias and data augmentation
@@
Deep neural networks are prone to suffer from data bias (Torralba and Efros, 2011; Fabbrizzi et al., 2022) and overfitting problems (Bejani and Ghatee, 2021). Overfitting can lead to the inability of trained models to generalize in real-world applications (Kornblith et al., 2019). Although overfitting can sometimes be detected by inspecting the training and validation performance over the course of training, it can often be imperceivable in challenging problems. A useful technique to detect overfitting is to use adversarial examples (Werpachowski et al., 2019) generated from the original dataset. On the other hand, it is important to address overfitting when the amount of data is limited. Data augmentation (Shorten and Khoshgoftaar, 2019; Rebuffi et al., 2021) techniques are usually employed to improve the robustness of deep neural networks.
2.4. Metamerism
@@
Metamerism is a physical phenomenon where distinct spectra produce the same color (Akbarinia and Gegenfurtner, 2018) as the high-dimensional spectral space is projected down to three dimensions of a trichromatic vision system (either the human eye or an RGB camera). This phenomenon has been studied in color science (Thornton, 1998; Foster et al., 2006), spectral rendering (Jakob and Hanika, 2019; Weidlich et al., 2021; Van De Ruit and Eisemann, 2023), and hyperspectral imaging (Hill, 1999; Foster and Amano, 2019). In hyperspectral imaging (Cukierski and Foran, 2010; Ortega et al., 2020), it is crucial for many applications to distinguish between metamers or near-metamers (i.e., different spectra that project to similar RGB values). Indeed spectral imaging is usually employed when the RGB color differences between two materials or features are too small to reliably distinguish between them. Therefore, hyperspectral imaging systems require special attention in the system design to acquire accurate spectral signatures (Hill, 2002; Ji et al., 2023).
3. Fundamentals
3.1. Spectral image formation
@@
We denote the hyperspectral image as a matrix ${\mathbf{X}}\in\mathbb{R}^{MN\times K}$, where $M,N$ are the number of pixels, and $K$ is the number of spectral bands. Note that we have stacked the 2D spatial dimensions in rows of ${\mathbf{X}}$. The spectral response function (SRF) of the camera can be expressed as a matrix ${\mathbf{Q}}\in\mathbb{R}^{K\times 3}$. Therefore, the spectrum-to-color projection results in a color image
(1) ${\mathbf{Y}}={\mathbf{X}}{\mathbf{Q}},$
@@
where ${\mathbf{Y}}\in\mathbb{R}^{MN\times 3}$. This is the color formation model in the NTIRE 2022 challenge (Arad et al., 2022). The inverse problem is to recover ${\mathbf{X}}$ from ${\mathbf{Y}}$.
@@
In the past NTIRE challenges (Arad et al., 2018, 2020, 2022), optical aberrations haven’t been included in the image simulation pipeline. However, the optical system of the RGB camera inevitably introduces spectrally-varying blurs to the spectral images, which is modeled as PSFs. This optical process can be described by a linear matrix-vector product in each spectral band followed by a sum over the spectral dimension. The spectral images through the optical system are ${\mathbf{W}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right)$, where ${\mathbf{A}}\in\mathbb{R}^{KMN\times MN}$ is a block matrix that stacks the spectral PSF matrices vertically, and $\mathrm{diag}(\cdot)$ extracts the diagonal blocks. The final RGB image is then
(2) ${\mathbf{Z}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right){\mathbf{Q}},$
@@
where ${\mathbf{Z}}\in\mathbb{R}^{MN\times 3}$. With the optical image formation model accounted for, the inverse problem is to recover ${\mathbf{X}}$ from ${\mathbf{Z}}$. See Appendix for the detailed image formation models.
@@
3.2. Hyperspectral datasets and data diversity
Compared to very large color (RGB) image datasets (e.g., ImageNet (Deng et al., 2009), DIV2K (Agustsson and Timofte, 2017)), hyperspectral datasets are far smaller in size, primarily limited by the unavailability of high-quality hyperspectral cameras and the difficulty in acquiring outside the lab with moving target scenes. The largest dataset so far is ARAD1K used in the NTIRE 2022 challenge (Arad et al., 2022). In addition, we also include the CAVE (Yasuma et al., 2010), ICVL (Arad and Ben-Shahar, 2016), and KAUST (Li et al., 2021) datasets that share the same spectral range to extend our experiments. The datasets are summarized in Table 1. Although other datasets (Chakrabarti and Zickler, 2011; Choi et al., 2017a; Monno et al., 2015) exist, they cover slightly different spectral bands, making direct comparisons difficult. We therefore restrict our analysis to the datasets listed in the table.
Table 1. Basic information of four existing hyperspectral datasets.
@@ 3.2. Hyperspectral datasets and data diversity
CAVE (Yasuma et al., 2010) 400:10:700 512 $\times$ 512 $\times$ 31 32 monochrome sensor + tunable filters lab setup
ICVL (Arad and Ben-Shahar, 2016) 400:10:700 1392 $\times$ 1300 $\times$ 31 201 HS camera (Specim PS Kappa DX4) outdoor
KAUST (Li et al., 2021) 400:10:700 512 $\times$ 512 $\times$ 31 409 HS camera (Specim IQ) outdoor
ARAD1K (Arad et al., 2022) 400:10:700 482 $\times$ 512 $\times$ 31 1000 HS camera (Specim IQ) outdoor
The difficulties in the data capture not only affect the size but also the diversity of the datasets. In particular, effects like metamerism, which are comparatively rare in everyday environments (Foster et al., 2006; Prasad and Wenhe, 2015), yet crucial for many actual applications of spectral imaging, are under-represented in the datasets. While the CAVE dataset (Yasuma et al., 2010) contains some fake-and-real pairs of objects (e.g., Fig. 2) to account for metamerism, the total amount of such data is still very low. We analyze the general data diversity issue in Section 4 and the specific case of metamerism in Section 5.
3.3. Modeling metamerism
Since there are not enough examples of metamerism in existing datasets, their effects in spectral reconstruction went unnoticed in prior works. On the one hand, we need adversarial examples to reveal the unexplored problems. On the other hand, we want to investigate how they can complement the current datasets. Therefore, we propose a new form of data augmentation in our experiments. Metameric augmentation starts with existing spectral images and creates a new, different spectral image that however maps to the same RGB image (given a specific set of RGB spectral response functions). In Section 5, we first use metameric augmentation as an adversarial example to reveal the previously omitted effects of metamerism on the performance discrepancy. In Section 6, we show metameric augmentation is beneficial to mitigate the raised problems along with an aberration-aware training strategy.

### v3

# Limitations of Data-Driven Spectral Reconstruction: Optics-Aware Analysis and Mitigation

@@ Abstract
I Introduction
@@
Hyperspectral imaging is a method that involves recording the light in a scene in the form of many, relatively narrow, spectral bands, rather than projected into three broadband RGB color channels. Where RGB imaging utilizes the trichromaticity theory of human color vision, spectral imaging provides additional information that can help discriminate between different materials and lighting conditions that are hard to tell apart in RGB images. For example, red stains in a crime scene could be blood, or paint, or a dyed cloth, which cannot be distinguished from their RGB colors. Skin tumors could not be diagnosed from surrounding tissues of the same color. It is difficult to spot and sort out plastic leaves from living plants by their greenish colors. Therefore, spectral imaging has been applied in many fields, including computer graphics [38], machine vision [25, 51], healthcare [45], agriculture [22], and environment [10], to name just a few.
@@
However, conventional hyperspectral cameras require scanning mechanisms [44, 32] to acquire the 3D hyperspectral datacube with 2D sensors. To simplify the demanding hardware, extensive efforts have been made in the development of various snapshot hyperspectral cameras [62, 53, 36].
@@
On the extreme end of these hardware simplification efforts, deep learning methods have emerged in recent years that attempt to solve the problem entirely in software by reconstructing spectral data from RGB images (RGB2HS). This has resulted in three CVPR-hosted NTIRE challenges [5, 6, 7] and various network architectures [54, 41, 72, 14, 71]. Yet it remains unclear how these methods generalize to unseen data, how they deal with the difficult but important problem of resolving metamerism [29, 47], and how they depend on the optical system of both the RGB source and the spectral cameras used to capture the datasets.
@@
At its core, estimating spectral information from RGB colors is an under-determined one-to-many mapping problem. As stated By Pharr et al. [49] (Ch. 4), “any such conversion is inherently ambiguous due to the existence of metamers”. Intuition would therefore suggest that the achievable spectral fidelity of RGB to spectral methods is limited. However, this intuition flies in the face of very high numerical test results reported in recent NTIRE challenges [5, 6, 7].
One possible explanation for the experimental success of spectral reconstruction from an RGB image is that the networks learn to exploit spatial structures, or scene semantics, to estimate spectral information. However, spectral images are usually employed when RGB images do not provide sufficient information for downstream tasks. Therefore, it is questionable to use scene semantics to resolve spectral ambiguities. Instead, in computational imaging, the idea is to measure the spectral information to help better understand the scene semantics, particularly in difficult scenarios. Clearly, these two methodologies feature reversed information flow. We argue that the computational imaging approach is more compatible with spectral imaging itself.
@@
In practice, metameric or near-metameric colors often occur in situations where the spatial structure is also similar, e.g., vein finding (the global geometry is a segment of a forearm, but vein structure is unknown), or biometrics (e.g., distinguishing real faces from masks or images). To illustrate this effect, we show an example from the smaller and older CAVE dataset [67] where several fake and real objects are presented in two groups, as shown in Fig. 1. Two sample points from the two red peppers (one real and the other fake, but it’s unclear which is which) have the RGB values (86, 21, 10) and (86, 21, 8), respectively. As can be seen, the real spectra of the real and fake red peppers differ substantially in the red part of the spectrum. A pre-trained model (arbitrarily chosen from previous NTIRE challenges trained on the much larger and newer ARAD1K dataset [7]) predicts spectra that differ from the ground truth, but more importantly, the predicted spectra for the real and fake pepper are almost identical, illustrating the failure to adequately deal with metamers.
@@
Figure 1: An example scene fake_and_real_food_ms from the CAVE dataset [67] consists of objects with visually similar colors, but actually different spectra. Left: Color image with highlighted points on the red peppers. Their RGB values are nearly the same. Right: Ground-truth and reconstructed spectra at the corresponding points show their spectral differences. The reconstructed spectra are predicted by the pre-trained MST++ model [14] on the ARAD1K dataset [7]. The neural network struggles to distinguish either the two spectra from each other, or from their true spectra.
@@
  • Crucially, the limitations of the datasets that we document not only affect the RGB to spectral work, but also any other spectral reconstruction and processing that uses the same training data [34, 65, 8]. We show how metameric augmentation can be used to at least partially overcome the dataset issues.
@@
A seemingly apparent observation from the results we show in this paper reinforces that it is impossible to distinguish metamers solely from RGB colors. Remarkably, this fundamental limitation has been largely overlooked within the research in this field. Through the evidence in this work, we contribute to a deepened understanding of the limitations of current datasets as well as of underlying sources that result in the limitations of spectral reconstruction accuracy. The results of the interplay between metameric spectra and optical aberrations open the door for new approaches for spectral recovery down the road.
II Related Work
II-A Hyperspectral cameras
@@
Conventional hyperspectral imaging systems require filter wheels, liquid-crystal tunable filters, or mechanical motion (e.g., pushbroom) [44, 32] to scan the 3D hyperspectral datacube. To enable snapshot acquisition, coded-aperture snapshot spectral imager (CASSI) [62, 19] has been proposed to achieve high spectral accuracy using spectrum-dependent coded patterns. Recent methods also exploit spectrally encoded point spread functions (PSFs) to computationally reconstruct a hyperspectral image [36, 9, 15]. In general, great efforts have been made to simplify hyperspectral camera hardware by software reconstruction.
II-B Spectral reconstruction from RGB images
@@
A recent trend to solve the snapshot hyperspectral imaging problem is to exploit hyperspectral data with deep neural networks to reconstruct spectral information from RGB images [4]. Owing to the wide availability of RGB cameras, this approach seems to be a promising candidate for hyperspectral imaging if successful. A large number of neural network architectures have been proposed in the past three NTIRE spectral recovery challenges  [5, 6, 7]. Our analysis in this paper focuses on this class of methods to gain insights on their strengths and limitations.
II-C Dataset bias and data augmentation
@@
Deep neural networks are prone to suffer from data bias [57, 24] and overfitting problems [11]. Overfitting can lead to the inability of trained models to generalize in real-world applications [39]. Although overfitting can sometimes be detected by inspecting the training and validation performance over the course of training, it can often be imperceivable in challenging problems. A useful technique to detect overfitting is to use adversarial examples [64] generated from the original dataset. On the other hand, it is important to address overfitting when the amount of data is limited. Data augmentation [55, 52] techniques are usually employed to improve the robustness of deep neural networks.
II-D Metamerism
@@
Metamerism is a physical phenomenon where distinct spectra produce the same color [2] as the high-dimensional spectral space is projected down to three dimensions of a trichromatic vision system (either the human eye or an RGB camera). This phenomenon has been studied in color science [56, 28], spectral rendering [35, 63, 58], and hyperspectral imaging [29, 27]. In hyperspectral imaging [21, 47], it is crucial for many applications to distinguish between metamers or near-metamers (i.e., different spectra that project to similar RGB values). Indeed spectral imaging is usually employed when the RGB color differences between two materials or features are too small to reliably distinguish between them. Therefore, hyperspectral imaging systems require special attention in the system design to acquire accurate spectral signatures [30, 37].
III Fundamentals
Figure 2: Spectral image formation models used in the analysis in this work. Top: In the NTIRE spectral recovery challenges, an RGB image is considered as a linear projection from a high-dimensional hyperspectral datacube to a 3D color image. The existence of metamerism results in identical RGB images for different spectra. The neural network trained in this way cannot distinguish their corresponding spectra. Bottom: A possible mitigation to this problem is to include the optical aberrations of the lens in the image formation model. Spectral information is encoded into the aberrated RGB images, enabling the neural network to tell the difference between metamers. In both cases, the RGB image differences are shown on the right (intensity enhanced for better visualization).
III-A Spectral image formation
@@
We denote the hyperspectral image as a matrix ${\mathbf{X}}\in\mathbb{R}^{MN\times K}$, where $M,N$ are the number of pixels, and $K$ is the number of spectral bands. Note that we model the spectral radiance here, not spectral reflectance. Illumination spectrum is included. We have stacked the 2D spatial dimensions in rows of ${\mathbf{X}}$. The spectral response function (SRF) of the camera can be expressed as a matrix ${\mathbf{Q}}\in\mathbb{R}^{K\times 3}$. Therefore, the spectrum-to-color projection results in a color image
${\mathbf{Y}}={\mathbf{X}}{\mathbf{Q}},$ (1)
@@
where ${\mathbf{Y}}\in\mathbb{R}^{MN\times 3}$. This is the color formation model in the NTIRE 2022 challenge [7]. The inverse problem is to recover ${\mathbf{X}}$ from ${\mathbf{Y}}$.
@@
In the past NTIRE challenges [5, 6, 7], optical aberrations haven’t been included in the image simulation pipeline. However, the optical system of the RGB camera inevitably introduces spectrally-varying blurs to the spectral images, which is modeled as PSFs. This optical process can be described by a linear matrix-vector product in each spectral band followed by a sum over the spectral dimension. The spectral images through the optical system are ${\mathbf{W}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right)$, where ${\mathbf{A}}\in\mathbb{R}^{KMN\times MN}$ is a block matrix that stacks the spectral PSF matrices vertically, and $\mathrm{diag}(\cdot)$ extracts the diagonal blocks. The final RGB image is then
${\mathbf{Z}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right){\mathbf{Q}},$ (2)
@@
where ${\mathbf{Z}}\in\mathbb{R}^{MN\times 3}$. With the optical image formation model accounted for, the inverse problem is to recover ${\mathbf{X}}$ from ${\mathbf{Z}}$. It is evident that the optical property in ${\mathbf{A}}$ spreads the spectral information to the RGB channels, offering side-channel information to help spectral reconstruction. See the Appendix for the full derivation.
@@
III-B Hyperspectral datasets and data diversity
Compared to very large color (RGB) image datasets (e.g., ImageNet [23], DIV2K [1]), hyperspectral datasets are far smaller in size, primarily limited by the unavailability of high-quality hyperspectral cameras and the difficulty in acquiring outside the lab with moving target scenes. The largest dataset so far is ARAD1K used in the NTIRE 2022 challenge [7]. In addition, we also include the CAVE [67], ICVL [4], and KAUST [42] datasets that share the same spectral range to extend our experiments. The datasets are summarized in Table I. Although other datasets [16, 18, 46] exist, they cover slightly different spectral bands, making direct comparisons difficult. We therefore restrict our analysis to the datasets listed in the table.
TABLE I: Basic information of four existing hyperspectral datasets.
@@ III-B Hyperspectral datasets and data diversity
CAVE [67] 400:10:700 512 $\times$ 512 $\times$ 31 32 monochrome sensor + tunable filters lab setup
ICVL [4] 400:10:700 1392 $\times$ 1300 $\times$ 31 201 HS camera (Specim PS Kappa DX4) outdoor
KAUST [42] 400:10:700 512 $\times$ 512 $\times$ 31 409 HS camera (Specim IQ) outdoor
ARAD1K [7] 400:10:700 482 $\times$ 512 $\times$ 31 1000 HS camera (Specim IQ) outdoor
The difficulties in the data capture not only affect the size but also the diversity of the datasets. In particular, effects like metamerism, which are comparatively rare in everyday environments [28, 50], yet crucial for many actual applications of spectral imaging, are under-represented in the datasets. While the CAVE dataset [67] contains some fake-and-real pairs of objects (e.g., Fig. 1) to account for metamerism, the total amount of such data is still very low. We analyze the general data diversity issue in Section IV and the specific case of metamerism in Section V.
III-C Modeling metamerism
Since there are not enough examples of metamerism in existing datasets, their effects in spectral reconstruction went unnoticed in prior works. On the one hand, we need adversarial examples to reveal the unexplored problems. On the other hand, we want to investigate how they can complement the current datasets. Therefore, we propose a new form of data augmentation in our experiments. Metameric augmentation starts with existing spectral images and creates a new, different spectral image that however maps to the same RGB image (given a specific set of RGB spectral response functions). In Section V, we first use metameric augmentation as an adversarial example to reveal the previously omitted effects of metamerism on the performance discrepancy. In Section VI, we show metameric augmentation is beneficial to mitigate the raised problems along with an aberration-aware training strategy.